OpenStudy (anonymous):

What is 490% written as a decimal?

OpenStudy (amoodarya):

\[=\frac{ 490 }{ 100}\]

OpenStudy (anonymous):

correct ^

OpenStudy (amoodarya):

now simplify to simplest form

OpenStudy (anonymous):

.49

OpenStudy (amoodarya):

\[\frac{ 490 }{100}=\frac{ 49 }{10}=4.9\]

OpenStudy (anonymous):

To make a percent into a decimal, move the decimal dot two places to the left. In this case 490% = 4.90 = 4.9

OpenStudy (anonymous):

I didnt want to give the answer , where did he go ?